OpenStudy (anonymous):

\[\log_{7} 343+\log_{2} \sqrt{32}-\log_{1/2} (1/2)\]

OpenStudy (anonymous):

\[\LARGE \log_{n}{x} = \frac{\log{x}}{\log{n}}\] \[\Large \log_{7}{343} +\log_{2}{\sqrt{32}} -\log_{(1/2)}{1/2} = \frac{\log{343}}{\log{7}}+\frac{\log{\sqrt{32}}}{\log{2}}-\frac{\log{1/2}}{\log{1/2}}\]

OpenStudy (anonymous):

write all those numbers, inside the logarithms, as the base of the logarithm, raised to some power. Then it's pretty easy to work out what it comes out as.

OpenStudy (anonymous):

the answer is 9/2

OpenStudy (anonymous):

@nitz please don't just give out the answer. That's against the code of conduct.

OpenStudy (anonymous):

\[\log(7^{3})/\log7+\log \sqrt{2^{5}}/\log 2-\log(1/2)/\log(1/2)\]

OpenStudy (anonymous):

I think they want you to use factorization on those inside numbers. So for example 343 = 7^3. Then, log_{a} {b} essentially is asking to solve a^x = b for x. So for that first term, you need to solve 7^x = 343. And I just said 7^3 = 343. Also, for the second term you can simplify that a bit and put it as 2^x = sqrt(32)

OpenStudy (apoorvk):

7^3 = 343 and \(\log_nM^a = a\log_nM\) use these two fundas to solve^.

OpenStudy (anonymous):

now you can use \[\log(m ^{n})=nlog(m)\]

OpenStudy (anonymous):

\[\log_{7} 343=x 7^{x} = 7^{3} x=3\] \[\log_{2}\sqrt{343} =x 2^{x} =2^{5} x=5\] \[\log_{1/2} (1/2)=x 1/2^{x} = 1/2 x=1/2\] 3+5-1/2=7/2 is it right?? @nitz @apoorvk @scarydoor

OpenStudy (anonymous):

first line yes, next two lines, no.

OpenStudy (anonymous):

actually I don't know what you're doing with those x's, but the first and last part of the first line are right.

OpenStudy (anonymous):

sorry is \[\sqrt{32}\] not 343

OpenStudy (anonymous):

\[\log(\sqrt{32})/\log(2)=\log(\sqrt{2^{5}})/(\log(2))=(5/2)*(\log 2)/(\log2)=5/2\]

OpenStudy (anonymous):

32 = 2*16 = 2*2^2. So sqrt(32) = 2*sqrt(2). log_2 (sqrt(32)) is equivalent to asking 2^x = sqrt(32). and sqrt(32) = 2*sqrt(2) = 2^(3/2). log_(1/2) (1/2) = x is equivalent to asking (1/2) ^ x = (1/2). x is clearly equal to 1.

OpenStudy (anonymous):

\[\log_{1/2} (1/2)\] how to solve @nitz

OpenStudy (anonymous):

sorry my first line is wrong. 16 = 4^2 = 2^4 ... so adjust the numbers for that.

OpenStudy (anonymous):

\[\log _{n}(m)=\log(m)/\log(n)\] \[\log(1/2)/\log(1/2)=1\]

OpenStudy (anonymous):

I suggest just remembering what the logarithm function was invented for. You had exponential functions where you have a^x = y, for some base a. Then people wanted to take the inverse of that. What I mean is that if they knew that a^x = y and they knew both a and y, they wanted a function that worked the other way. This function is log_a (y) = x. You just need to figure out "a to the power [something] equals y". since you can factor all the y's in terms of a, then it's easy to work out and you don't have to bother with any kind of change of base. I think that actually makes it harder in this case.

OpenStudy (anonymous):

and think about "half to the power [something] is equal to a half" what is the "something"? it clearly has to be one. Messing around with it any more is just making it more complicated for yourself I think.

OpenStudy (anonymous):

sometimes the notation of the logarithm function confuses people away from the simplicity of what it's asking...

OpenStudy (anonymous):

ln and log is the same ?

OpenStudy (anonymous):

I think there are different standards. Where I am, ln is used when the base is the number e. ln is then called the natural logarithm, because e is used a lot in "natural" things / science. ln x = log_e x.

OpenStudy (anonymous):

ln-natural log with base e log -with base 10

OpenStudy (anonymous):

occasionally log is used without mentioning the base when the base is 2. It depends on the people who are using the maths usually... computer scientists often use log to mean base 2 unless otherwise specified. Probably base 10 is more common though, if the base isn't specified.

OpenStudy (anonymous):

k thanks :) All @nitz @scarydoor @Wired @apoorvk

OpenStudy (anonymous):

for example this quest ln x= ln17+ ln13 how to solve with ln

OpenStudy (apoorvk):

Use the funda --> \(log_a m + \log_a n = =\log_a(mn)\)

OpenStudy (apoorvk):

*single '=' sign. And yes, then you remove the log signs (since bases are same).

OpenStudy (anonymous):

e is irrational. So, I'm pretty sure (haven't done this stuff in a while) that you can only solve the right hand side by using a calculator, or, at if you for some reason are told not to, then by a numerical method that will approximate it to some desired accuracy, which is what the calculator would essentially do. But the teacher wouldn't be asking that... So you solve right hand side. Say ln x = y, where you know what y is. Then raise e to the power of both sides, so you have e^(ln x) = x = e ^ y. Then you again have to use the calculator to solve e ^ y, and then you've solved for x.

OpenStudy (anonymous):

ln(17+13)=30 is it right?

OpenStudy (anonymous):

apoorvk's response above is better than mine. I forgot you can combine them and then not bother about calculating any logarithms.

OpenStudy (maheshmeghwal9):

\[\log_{7} 343+\log_{2} \sqrt{32}-\log_{1/2} (1/2)\]\[\implies \log_7{7^3} +\log_2{2^{5*\frac{1}{2}}} -\log_{2^{-1}}{2^{-1}}.\]\[\implies 3\log_7{7}+5*\frac{1}{2}\log_2{2}-(\frac{-1}{-1})\log_2{2} .\]\[\implies 3+\frac{5}{2}-1=?\]It is becoz \[\log_a{a}=1.\]

OpenStudy (maheshmeghwal9):

do this urself now:)

OpenStudy (anonymous):

by what apoorvk said above, ln 17 + ln 13 = ln(17*13). Then left hand side and right hand side are the same number. So take the number e and raise that by the number on left hand side, and that's equal to e raised to the number on right hand side. Then remember that e ^ ln(x) = x. To see why that's true, remember what the ln function does. it's an inverse of e^x function. So then you have x = 17 * 13. I think. Unless I said something silly.

OpenStudy (farmdawgnation):

@nitz Please do refrain from giving out answers. Thanks.